Hydrogen is placed along with Alkali metals in the modern periodic table though it shows non-metallic character a) as Hydrogen has one electron & readily loses electron to form negative ion b) as Hydrogen can easily lose one electron like alkali metals to form positive ion c) as Hydrogen can gain one electron easily like Halogens to form negative ion d) as Hydrogen shows the properties of non-metals

Grade:10

11 Answers

Gaurav Yadav
3 years ago
Hey student.
 
b) As Hydrogen can easily lose one electron like alkali metals to form positive ion
 
This is the answer to your question. Hope this helps you.
Nandini Sharma
3 years ago
Which of the following has highest electronegativity? a) F b) Cl c) Br d) I
Which of the following has highest electronegativity? a) F b) Cl c) Br d) I
Gaurav Yadav
3 years ago
Hey student.
 
a) F 
 
This is the answer to your question. Hope this helps you.
This is the answer to your question. Hope this helps you.
Nandini Sharma
3 years ago
Identify the reason for the gradual change in electronegativity in halogens down the group. a) Electronegativity increases down the group due to decrease in atomic size b) Electronegativity decreases down the group due to decrease in tendency to lose electrons c) Electronegativity decreases down the group due to increase in atomic radius/ tendency to gain electron decreases d) Electronegativity increases down the group due to increase in forces of attractions between nucleus & valence electrons
Gaurav Yadav
3 years ago
Hey student.
 
c) Electronegativity decreases down the group due to increase in atomic radius/ tendency to gain electron decreases.
 
This is the answer to your question. Hope this helps you.
Nandini Sharma
3 years ago
Which of the following reason correctly justifies that “Fluorine (72pm) has smaller atomic radius than Lithium (152pm)”? a) F and Li are in the same group. Atomic size increases down the group b) F and Li are in the same period. Atomic size increases across the period due to increase in number of shells c) F and Li are in the same group. Atomic size decreases down the group d) F and Li are in the same period and across the period atomic size/radius decreases from left to right.
 
Gaurav Yadav
3 years ago
Hey student.
 
d) F and Li are in the same period and across the period atomic size/radius decreases from left to right.
 
This is the answer to your question. Hope this helps you.
Nandini Sharma
3 years ago
If the powers of the lenses L1 and L2 are in the ratio of 4:1, what would be the ratio of the focal length of L1 and L2? a) 4:1 b) 1:4 c) 2:1 d) 1:1
Gaurav Yadav
3 years ago
Hey student.
 
P=1/f P1=1/f1 and P2=1/f2 P1/P2=4/1, hence (1/f1)/(1/f2) = 4/1 Hence f1/f2 = ¼ 
b) ¼
 
This is the answer to your question. Hope this helps you.
Nandini Sharma
3 years ago
What is the formula for magnification obtained with a lens? a) Ratio of height of image to height of object b) Double the focal length. c) Inverse of the radius of curvature. d) Inverse of the object distance.
Gaurav Yadav
3 years ago
Hey student.
 
a) Ratio of height of image to height of object
 
This is the answer to your question. Hope this helps you.